Consider a hypothesis economy described by the following equation

C=100

I=1200

X-1110

M=200+0.25Y

T=250+0.3Y

Required

1.Compute the equilibrium level of the national income

2.The level of consumption income after tax and net exports that corresponds to the equilibrium level

national income


1
Expert's answer
2021-06-25T11:06:33-0400

1.

"Y=C+I+G+X-M"

"Y=100+1200+0+1110-(200+0.25Y)"

"Y=1300+1110-200-0.25Y"

"Y+0.25Y=2410-200"

"1.25Y=2210"

"Y=\frac{2210}{1.25}"

National income Y = 1768

2.

Level of consumption income after tax "(Y_d)=Y-Tax(T)"

"Y=1768"

"T=250+0.3Y=250+0.3(1768)"

"T=780.4"

"Y_D=Y-T=1768-780.4"

"Y_D=987.6"


Net exports=X-M

"X=1110"

"M=200+0.25Y"

"M=200+0.25(1768)"

"M=200+442"

"M=642"


Net exports "=1110-642=468"

The level of consumption income after tax is 987.6 and net exports is 468.